OpenStudy (anonymous):

Transform the following equations into standard form. 4x+3=x^2

OpenStudy (anonymous):

x^2-4x-3=0 -4+/-(sqrt(16+12)/2 -4+/-2rad7/2 -2+/-rad7 x=-2+RAD(7) x=-2-RAD(7)

OpenStudy (anonymous):

i dont knoww hat your asking so i just solved the equation